OpenStudy (anonymous):

how many three-digits integers can be divided y 4 to produce a new integer with the same ten' digit as the original integer?

OpenStudy (anonymous):

Maybe it will help to know that there are 225 numbers from 100-999 that are divisible by 4? 4*25 = 100 4*249 = 996

OpenStudy (anonymous):

not that i know a concrete way to figure this out, just throwing stuff out.

OpenStudy (anonymous):

which numbers?

OpenStudy (anonymous):

128 /4 = 32 those dont have the same tens digit.

OpenStudy (anonymous):

512/4 = 128, those also dont have the same tens digit.

OpenStudy (anonymous):

136/4 = 34, thats a match. So is 132/4 = 33

OpenStudy (anonymous):

hey joemath come to twiddla we r gonna find a real method to solve it

OpenStudy (anonymous):

http://www.twiddla.com/584034

OpenStudy (anonymous):

im down, im very curious as to a method other than guess and check lol

OpenStudy (anonymous):

We know that , N = Dq + R According to question, 100x+10y+z = 4(10y +z) +0 => 100x = 30y + 3 z = 3(10y+z) clearly Multiple of 100

OpenStudy (anonymous):

right

OpenStudy (anonymous):

like 300,303...

OpenStudy (anonymous):

no

OpenStudy (anonymous):

let me think again

OpenStudy (anonymous):

i think you are assuming the numbers will have the same ones digit. 100x+10y+z is the number xyz and 10y+z is the number yz

OpenStudy (anonymous):

yes .that is wrong.it should have same tens digit

OpenStudy (anonymous):

do we have more examples ...

OpenStudy (anonymous):

for the problem

OpenStudy (anonymous):

now 132 is one because 4*3 =12 and the first two term is 13 similarly in 200s 260 as 4*6 =24 and first two terms is 26

OpenStudy (anonymous):

nvm, im retarded

OpenStudy (anonymous):

and sleepy. i found some more numbers though. 260/4 = 65 264/4 = 66 268/4 = 67

OpenStudy (anonymous):

im sorta thinking about it like this: counting by ones, it takes 10 ones digit to change the tens digit. counting by 4's, it about 2 or 3 multiples to change the tens digit. think of this problem like "one guy completes a lap in 10 mins, the other one completes a lap in 2-3 mins, how many times with they meet on the track"

OpenStudy (anonymous):

the first time we hit a 3 digit number that satisfies this condition is 132: 132/4 = 33 the LCM of 10 and 3 is 30, so approximately around 33+30 i should find the next solution: 63* 4 = 252 nope 64*4 = 256 nope 65*4 = 260 bingo 66*4 = 264 bingo 67*4 = 268 bingo

OpenStudy (anonymous):

Adding another 30 to get in the general area again: 67+30 = 97 97*4 = 388 nope 98*4 = 392 bingo 99*4 = 396 bingo Do you guys see it yet?

OpenStudy (anonymous):

3 digit numbers divisible by 4 have their last 2 digits divisible by 4.

OpenStudy (anonymous):

I think there should 12 integers like this upto 408

OpenStudy (anonymous):

see thats the thing I was talking about

OpenStudy (anonymous):

4*3 =12 and we need some three digit number then as tens should be same ...which is only possible for 130s then we go to 4*4 =16 but not possible then 4*5 =20 which is not possible again but when we go into 4*6 =24 We get possibilities in 260s Now in every whether its 130s or 260s only 3 cases are possible then similar goes until 390 as 3*9 =36..and now 400 /4 =100 here also 3 cases are possible with totaling 12 upto 410